OpenStudy (anonymous):

PLEASE HELP! Given x/x^2-9 - x/x^2+6x+9, perform the indicated operation and simplify the result. Use your work to answer the following question. What is the LCD?

OpenStudy (lgbasallote):

look at your two denoms.. the first denom is x^2-9 = (x+3)(x-3) second denom is x^2 + 6x +9 = (x+3)^2 what do you think is the LCD?

OpenStudy (anonymous):

(x-3)(x+3)^2?